# Edge Zeta Functions and Eigenvalues for Spherical Buildings: A Uniform Hecke Algebra Approach

### v2

@@ Introduction
Zeta functions of spherical buildings reveal deep connections between geometry, combinatorics, and representation theory. They generalize the Ihara zeta functions of regular graphs [8] and their two–variable extensions for qq–regular trees [7]. In higher rank, the geometry of a building reflects the combinatorics of its Weyl group; for buildings arising from algebraic groups, it also encodes the underlying symmetries of the group, linking combinatorial and algebraic structure. The corresponding zeta function records weighted counts of closed geodesics or galleries, unifying spectral, combinatorial, and arithmetic data in a single analytic object.
The one–dimensional case, corresponding to regular graphs and hence to spherical buildings of dimension $1$, is completely understood through the classical theory of Ihara zeta functions and their variants [5, 15]. Motivated by this success, subsequent work extended the theory to higher–dimensional analogues arising from affine buildings of reductive groups over local fields. In particular, the edge and gallery zeta functions of finite quotients of two–dimensional affine buildings have been extensively investigated, revealing deep connections with pp–adic LL–functions and automorphic forms ([10, 13, 11, 4, 14]). By contrast, the zeta functions of spherical buildings over finite fields remain much less explored. Beyond the one–dimensional case, previous studies have focused mainly on oppositeness graphs and their spectra [1, 19], leaving the general geodesic structure largely unexplored.
In earlier work [18], the present author introduced the edge zeta function for spherical buildings of type $\mathbf{A}_{n}$ and derived closed formulas via a character–theoretic analysis of induced Weyl–group representations. This resolved the type $\mathbf{A}$ case completely and called for a generalization to other classical and exceptional types.
@@
Our main theorem 1.0.1 asserts that every nonzero edge eigenvalue $\lambda$ becomes a pure power of $q$ after being raised to an integer exponent $k=2m$ depending on the Lie type and the type orbit. This establishes a representation–theoretic explanation of the “power–of–qq” phenomenon first observed in type $\mathbf{A}$, and shows that the same structure persists across all finite types. The proof combines Luo’s recent decomposition theorem [16] with Springer’s theorem on central elements of Iwahori–Hecke algebras [6, §9.2]. A refinement of the main result (Theorem 3.3.1) further gives an explicit description of the associated spectral data and the zeta functions. Finally, we derive new closed formulas for the symplectic case (type $\mathbf{C}$), tabulate the corresponding cycle structures and the exponents $2m$.
@@
  • Section 4 presents the explicit computations—recovering type $\mathbf{A}$ formulas, deriving new formulas for type $\mathbf{C}$, and listing cycle data and half–lengths $m$ for all finite types (except type $\mathbf{D}$).
In this way, the paper extends the type $\mathbf{A}$ results of [18] to all finite spherical buildings arising from groups of Lie type, showing that the edge zeta functions are uniformly controlled by the representation theory of Hecke algebras across all finite Lie types.
@@ 1 Settings and Main Theorem
For type $\mathbf{A}$, one may take $k=6$ for all $\lambda$ ([18, §7.2, 8.4]). For types $\mathbf{B}$ or $\mathbf{C}$, one may take $k=8$ (Section 4.3), and for type $\mathbf{D}$, $k$ can be 66 or 88 depending on $\lambda$. Explicit $k$ (or $m=k/2$) values for various graphs appear in the Section $4.$
@@
For any linear operator $T$, write $\mathop{\rm Spec}\nolimits(T)=\{\text{multiset of eigenvalues of }T\}$. For a finite graph $X$, $\mathop{\rm Spec}\nolimits(X)$ denotes the multiset of eigenvalues of its adjacency operator.
Throughout, $e_{P}$ denotes the normalized idempotent $|P|^{-1}\sum_{x\in P}x$ in the group algebra $\mathbb{C}G$, and $a_{PwQ}$ denotes the double–coset sum $a_{PwQ}=|Q|^{-1}\sum_{x\in PwQ}x$. All tensor products are taken over $\mathbb{C}$.
@@ 2.1 From geodesic cycles to the geodesic edge graph
Remark. $X_{2}(\mathcal{B})$ was introduced in [18, §2] under the name geodesic edge graph. From a graph-theoretic perspective, $X_{2}(\mathcal{B})$ is a subgraph of the oriented line graph of $X_{1}(\mathcal{B})$, whose vertices represent directed edges of $X_{1}(\mathcal{B})$ and where adjacency encodes consecutive edge-pairs (see [12]). In our case, the adjacency relation is further restricted by the local geodesic condition of Lemma 2.1.1.
@@ 2.3 Transitive Partite Action and RDE Operators
Left multiplication by $G$ acts transitively on each type class and preserves adjacency (as per Lemma 2.1.1 (ii)). Thus, $G$ acts partite-transitively on $X_{2}(\mathcal{B})|_{C}$ in terms of [18, § 3.2].
@@
We now recall the group algebra method to find zeta functions for partite-transitive graphs(cf. [18, §2–3]):
@@
    where $\lambda$ runs over the eigenvalues of $T_{C}$ (See for example [18] Prop. 2.1.11 and Prop. 3.2.9). Hence, the nonzero eigenvalues of $X_{2}(\mathcal{B})|_{C}$ are the $c$-th roots of the eigenvalues of $T_{C}$.
    Remark. Each map $T_{i}$ is a Hecke operator: it acts by summing over the double coset $P_{i}w_{i+1}P_{i+1}$ determined by the local type transition $(r_{i}\to r_{i+1})$. Thus, $T_{C}$ represents a composite Hecke operator corresponding to the $c$-step product of such double cosets.
@@
    For $g\in G$, there is an edge from $v_{i}$ to $gv_{i+1}$ if and only if $gP_{i+1}\subseteq P_{i}w_{S-\{r_{i+1}\}}P_{i+1}$. Write the maximal length element $w_{S-\{r_{i+1}\}}$ in $W_{S-\{r_{i+1}\}}$ as $w_{i+1}$. Then we have
@@
    $D(v_{i},v_{i+1},1)=\alpha_{v_{i+1}}\left(\sum_{v_{i}\to w}w\right)=\sum_{gP_{i+1}\subseteq P_{i}w_{i+1}P_{i+1}}ge_{P_{i+1}}=a_{P_{i}w_{i+1}P_{i+1}},$
    where $a_{PwQ}$ denotes the element
    $a_{PwQ}=\frac{1}{|Q|}\sum_{x\in PxQ}x\quad(w\in W;P,Q\text{ are parabolic subgroups of }G).$
@@
    $Z_{C}\!\left(X_{2}(\mathcal{B}),u\right)=\prod_{\lambda}\frac{1}{1-\lambda u^{c}},$
@@
In the next subsection, we simplify $D(v_{0},v_{0},c)$ using Weyl group length-additivity and Luo’s theorem on reduced products, collapsing the above product to a small number of double-coset operators. This, together with Springer’s theorem (as in [1, 18]), yields that the resulting eigenvalues are powers of $q$, completing the proof of Theorem 1.0.1 componentwise.
@@ 3.1 Collapsing the RDE product
where $D(v_{i},v_{i+1},1)=a_{P_{i}w_{i+1}P_{i+1}}\in e_{P_{i}}\mathbb{C}Ge_{P_{i+1}}$, as discussed in §2.3. Here, and $w_{i}=w_{S-\{r_{i}\}}$ is the longest element in $W_{S-\{r_{i}\}}$.
@@
Let $w^{\prime}_{i}$ be the longest element in $W_{S-\{i,i+1\}}$, and $w_{i}$ be the longest element in $W_{S-\{i\}}$. Then there exists a unique integer $m$ such that
$w^{\prime}_{0}w_{S}=w^{\prime}_{0}w_{1}w^{\prime}_{1}w_{2}\cdots w^{\prime}_{m-1}w_{m},$
and
$l(w^{\prime}_{0}w_{S})=\sum_{k=1}^{m}l(w^{\prime}_{k-1}w_{k}).$
@@
Let $\{s_{i}\}_{i\in\mathbb{Z}}$ be the sequence from Luo’s construction, and write
$w_{i}:=w_{S-\{s_{i}\}}\quad\text{and}\quad w^{\prime}_{i}:=w_{S-\{s_{i},s_{i+1}\}}.$
For each starting index $k\geq 0$, suppose Luo’s factorization holds with length $m_{k}$:
@@
By (2), $\sum_{j=1}^{m_{k}}\ell\!\bigl(w^{\prime}_{k+j-1}w_{k+j}\bigr)=\ell(w^{\prime}_{k}w_{S}).$
@@
$|BsB|=q_{s}\,|B|,$
so $q_{s}$ is the Hecke parameter attached to $s$, and may vary between different conjugacy classes of simple reflections.
@@
$Q(w)\;:=\;\prod_{s\in S}q_{s}^{m_{s}(w)},$
where $m_{s}(w)$ is the number of occurrences of $s$ in any reduced expression of $w$. This is well-defined because the parameters $q_{s}$ are constant on conjugacy classes of simple reflections. Then
@@
Recall (see Lemma 8.2 in [5]) that
@@
Thus
@@
and moreover, $I_{i-1}=S-\{s_{i-1},s_{i}\}$ is conjugated to $I_{i}=S-\{s_{i},s_{i+1}\}$ via $u_{i}$: $w^{\prime}_{i-1}$ is the longest element in $W_{I_{i-1}}$ and fixes $I_{i-1}$, while $w_{i}=w_{S-\{s_{i}\}}$ conjugates $I_{i-1}$ to $I_{i}$:
$w_{i}^{-1}(w^{\prime}_{i-1})^{-1}I_{i-1}w^{\prime}_{i-1}w_{i}=w_{i}^{-1}I_{i-1}w_{i}=I_{i}.$
@@ 3.2 Eigenvalues and edge–zeta factors
This fact is closely related to the theory of oppositeness graphs, cf. [1], and generalizes Proposition 7.0.1 of [18]. Both proofs, ours and the one in [1], are ultimately based on Springer’s theorem on the action of central elements in Iwahori–Hecke algebras.
@@
where $\mathcal{R}$ runs over all conjugacy classes of reflections in the Weyl group $W$, and $\chi_{1}$ denotes the corresponding character of $W$ associated with $\chi$. Then the eigenvalues of $(a_{P_{I}wP_{J}}\,a_{P_{J}wP_{I}})_{r}$ are
$Q(w_{I})^{-2}\cdot q^{\,f_{\chi}},$
with multiplicity $d_{\chi}n_{\chi}$, as $\chi$ ranges over all irreducible representations of $G$ with $n_{\chi}\neq 0$.
@@
Now, $a_{BwB}^{2}$ lies in the center of the Hecke algebra $e_{B}\mathbb{C}Ge_{B}$ (by Springer’s theorem). Since $e_{P_{I}}\in e_{B}\mathbb{C}Ge_{B}$, we have
@@
Second, we change sides (Lemma 6.1.1 of [18]): the eigenvalues of $(a_{BwB}^{2})_{r}$ on $\mathbb{C}Ge_{P_{I}}$ coincide with those of $(a_{BwB}^{2})_{l}$ on $e_{P_{I}}\mathbb{C}G$.
@@
Finally, by Springer’s theorem (Theorem 9.2.2 in [6]), the central element $a_{BwB}^{2}\in e_{B}\mathbb{C}Ge_{B}$ acts on $e_{B}M_{\chi}$, and on $e_{P}M_{\chi}\subseteq e_{B}M_{\chi}$, as scalar multiplication by $q^{f_{\chi}}$. Hence the eigenvalues of $(a_{P_{I}wP_{J}}a_{P_{J}wP_{I}})_{r}$ are
@@ 3.3 Summary: proof and refinement of the main theorem
    $Z_{C}\!\left(X_{2}(\mathcal{B}),u\right)\;=\;\prod_{\lambda\in\mathop{\rm Spec}\nolimits(T_{C})}\frac{1}{1-\lambda\,u^{\,c}}=\;\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,\left(Q(w_{I})^{-2}\,q^{\,f_{\chi}}\right)^{1/d}\,u^{\,c}\bigr)^{\,m_{C,\chi}(\zeta)d_{\chi}}}.$
We have now proved the main theorem, in fact in the following refined form:
@@
    $Z\!\left(X_{2}(\mathcal{B})|_{C},u\right)=\;\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,\left(Q(w_{I})^{-2}\,q^{\,f_{\chi}}\right)^{1/d}\,u^{\,c}\bigr)^{\,m_{C,\chi}(\zeta)d_{\chi}}}.$
@@
$Z\!\left(X_{2}(\mathcal{B}),u\right)\;=\;\prod_{C}Z\!\left(X_{2}(\mathcal{B})|_{C},u\right),$
@@
By Tits’ Deformation Principle [18, §8.1], [6, §7.4,8.1], under $q\to 1$ we have $\mathcal{H}\to\mathbb{C}W$, $P_{0}\to W_{0}$, and $D(v_{0},v_{0},c)\to a_{W_{0}uW_{0}}=\tfrac{1}{|W_{0}|}\sum_{x\in W_{0}uW_{0}}x$. The module $e_{P_{I}}M_{\chi}$ deforms to $e_{W_{I}}M_{1,\chi}$, where $M_{1,\chi}$ is the $\mathbb{C}W$-module corresponding to $\chi$. Thus $m_{C,\chi}(\zeta)$ equals the multiplicity of $\zeta$ as an eigenvalue of $a_{W_{0}uW_{0}}$ on $e_{W_{0}}M_{1,\chi}$.

### v3

@@ Introduction
Zeta functions of spherical buildings reveal deep connections between geometry, combinatorics, and representation theory. They generalize the Ihara zeta functions of regular graphs [undefg] and their two–variable extensions for qq–regular trees [undeff]. In higher rank, the geometry of a building reflects the combinatorics of its Weyl group; for buildings arising from algebraic groups, it also encodes the underlying symmetries of the group, linking combinatorial and algebraic structure. The corresponding zeta function records weighted counts of closed geodesics or galleries, unifying spectral, combinatorial, and arithmetic data in a single analytic object.
The 1–dimensional spherical buildings, as well as finite quotients of trees (1–dimensional affine buildings), form regular graphs whose spectral and zeta–function properties are completely understood through the classical theory of Ihara zeta functions and their variants [undefq, undefd, undefn]. Motivated by this success, subsequent work extended the theory to higher–dimensional analogues arising from affine buildings of reductive groups over local fields. In particular, the edge and gallery zeta functions of finite quotients of two–dimensional affine buildings have been extensively investigated, revealing deep connections with pp–adic LL–functions and automorphic forms ([undefi, undefl, undefc, undefj, undefm]). By contrast, the zeta functions of spherical buildings over finite fields remain much less explored. Beyond the one–dimensional case, previous studies have focused mainly on oppositeness graphs and their spectra [undef, undefs], leaving the general geodesic structure largely unexplored.
In earlier work [undefr], the present author introduced the edge zeta function for spherical buildings of type $\mathbf{A}_{n}$ and derived closed formulas via a character–theoretic analysis of induced Weyl–group representations. This resolved the type $\mathbf{A}$ case completely at the level of the 1–skeleton, and called for a generalization to other classical and exceptional types.
@@
Our main theorem 1.0.1 asserts that every nonzero edge eigenvalue $\lambda$ becomes a pure power of $q$ after being raised to an integer exponent $k=2m$ depending on the Lie type and the type orbit. This establishes a representation–theoretic explanation of the “power–of–qq” phenomenon first observed in type $\mathbf{A}$, and shows that the same structure persists across all finite types. The proof combines Luo’s recent decomposition theorem [undefo] with Springer’s theorem on central elements of Iwahori–Hecke algebras [undefe, §9.2]. A refinement of the main result (Theorem 3.3.1) further gives an explicit description of the associated spectral data and the zeta functions. Finally, we derive new closed formulas for the symplectic case (type $\mathbf{C}$), and tabulate the corresponding cycle structures and the exponents $2m$.
@@
  • Section 4 presents the explicit computations—recovering type $\mathbf{A}$ formulas, deriving new formulas for type $\mathbf{C}$, and listing cycle data and half–lengths $m$ for all finite types other than type $\mathbf{D}$.
In this way, the paper extends the type $\mathbf{A}$ results of [undefr] to all finite spherical buildings arising from groups of Lie type, showing that the edge zeta functions are uniformly controlled by the representation theory of Hecke algebras across all finite Lie types.
@@ 1 Settings and Main Theorem
For type $\mathbf{A}$, one may take $k=6$ for all $\lambda$ ([undefr, §7.2, 8.4]). For types $\mathbf{B}$ or $\mathbf{C}$, one may take $k=8$ (Section 4.3), and for type $\mathbf{D}$, $k$ can be 66 or 88 depending on $\lambda$. Explicit $k$ (or $m=k/2$) values for various graphs appear in the Section $4.$
@@
For any linear operator $T$, write $\mathop{\rm Spec}\nolimits(T)$ for the multiset of eigenvalues of $T$. For a finite directed graph $X$, $\mathop{\rm Spec}\nolimits(X)$ denotes the multiset of eigenvalues of its adjacency operator.
Throughout, $e_{P}$ denotes the idempotent $|P|^{-1}\sum_{x\in P}x$ in the group algebra $\mathbb{C}G$, and $a_{PwQ}$ denotes the double–coset sum $a_{PwQ}=|Q|^{-1}\sum_{x\in PwQ}x$. All tensor products are taken over $\mathbb{C}$.
@@ 2.1 From geodesic cycles to the geodesic edge graph
Remark. $X_{2}(\mathcal{B})$ was introduced in [undefr, §2] under the name geodesic edge graph. From a graph-theoretic perspective, $X_{2}(\mathcal{B})$ is a subgraph of the oriented line graph of $X_{1}(\mathcal{B})$, whose vertices represent directed edges of $X_{1}(\mathcal{B})$ and where adjacency encodes consecutive edge-pairs (see [undefk]). In our case, the adjacency relation is further restricted by the local geodesic condition of Lemma 2.1.1.
@@ 2.3 Transitive Partite Action and RDE Operators
Left multiplication by $G$ acts transitively on each type class and preserves adjacency (as per Lemma 2.1.1 (ii)). Thus, $G$ acts partite-transitively on $X_{2}(\mathcal{B})|_{C}$ in terms of [undefr, § 3.2].
@@
We now recall the group algebra method to find zeta functions for partite-transitive graphs(cf. [undefr, §2–3]):
@@
    where $\lambda$ runs over the eigenvalues of $T_{C}$ (See for example [undefr] Prop. 2.1.11 and Prop. 3.2.9). Hence, the nonzero eigenvalues of $X_{2}(\mathcal{B})|_{C}$ are the $c$-th roots of the eigenvalues of $T_{C}$.
    Remark. Each map $T_{i}$ is a Hecke operator: it acts by summing over the $w$ such that there is an edge from $v$ to $w$. Thus, $T_{C}$ represents a composite Hecke operator corresponding to the $c$-step product of such double cosets.
@@
    For $g\in G$, there is an edge from $v_{i}$ to $gv_{i+1}$ if and only if $gP_{i+1}\subseteq P_{i}w_{S-\{r_{i+1}\}}P_{i+1}$, by lemma 2.1.1. Write the maximal length element $w_{S-\{r_{i+1}\}}$ in $W_{S-\{r_{i+1}\}}$ as $w_{i+1}$. Then we have
@@
    $D(v_{i},v_{i+1},1)=\alpha_{v_{i+1}}\left(\sum_{v_{i}\to w}w\right)=\sum_{gP_{i+1}\subseteq P_{i}w_{i+1}P_{i+1}}ge_{P_{i+1}}=\frac{1}{|P_{i+1}|}\sum_{x\in P_{i}w_{i+1}P_{i+1}}x=a_{P_{i}w_{i+1}P_{i+1}}.$
@@
    $Z_{c}\!\left(X_{2}(\mathcal{B})|_{C},u\right)=\prod_{\lambda}\frac{1}{1-\lambda u^{c}},$
@@
In the next subsection, we simplify $D(v_{0},v_{0},c)$ using Weyl group length-additivity and Luo’s theorem on reduced products, collapsing the above product to a small number of double-coset operators. This, together with Springer’s theorem (as in [undef, undefr]), yields that the resulting eigenvalues are powers of $q$, completing the proof of Theorem 1.0.1 componentwise.
@@ 3.1 Collapsing the RDE product
where $D(v_{i},v_{i+1},1)=a_{P_{i}w_{i+1}P_{i+1}}\in e_{P_{i}}\mathbb{C}Ge_{P_{i+1}}$, as discussed in §2.3. Here $w_{i}=w_{S-\{r_{i}\}}$ is the longest element in $W_{S-\{r_{i}\}}$.
@@
Let $w^{\prime}_{i}$ be the longest element in $W_{S-\{i,i+1\}}$, and $w_{i}$ be the longest element in $W_{S-\{i\}}$. Then there exists a unique integer $m$ such that $w^{\prime}_{0}w_{S}=w^{\prime}_{0}w_{1}w^{\prime}_{1}w_{2}\cdots w^{\prime}_{m-1}w_{m},$ and $l(w^{\prime}_{0}w_{S})=\sum_{k=1}^{m}l(w^{\prime}_{k-1}w_{k}).$
@@
Let $\{s_{i}\}_{i\in\mathbb{Z}}$ be the sequence from Luo’s construction, and set $w_{i}:=w_{S-\{s_{i}\}}$ and $w^{\prime}_{i}:=w_{S-\{s_{i},s_{i+1}\}}$. For each starting index $k\geq 0$, suppose Luo’s factorization holds with length $m_{k}$:
@@
By equation (2), $\sum_{j=1}^{m_{k}}\ell\!\bigl(w^{\prime}_{k+j-1}w_{k+j}\bigr)=\ell(w^{\prime}_{k}w_{S}).$
@@
$Q(s)=\frac{|BsB|}{|B|},$
so $Q(s)$ is the Hecke parameter attached to $s$, and may vary across different conjugacy classes of simple reflections.
@@
$Q(w)\;:=\;\prod_{s\in S}Q(s)^{m_{s}(w)},$
where $m_{s}(w)$ is the number of occurrences of $s$ in any reduced expression of $w$. This is well-defined because the parameters $Q(s)$ are constant on conjugacy classes of simple reflections. Then
@@
Recall (see Lemma 8.2 in [undefd]) that
@@
Since $w$ has minimal length in $W_{I}w$, we have $\ell(uw)=\ell(u)+\ell(w)$ and $Q(uw)=Q(u)Q(w)$ for any $u\in W_{I}$. It follows that
@@
and moreover, $I_{i-1}=S-\{s_{i-1},s_{i}\}$ is conjugated to $I_{i}=S-\{s_{i},s_{i+1}\}$ via $u_{i}$: $w^{\prime}_{i-1}$ is the longest element in $W_{I_{i-1}}$ and fixes $I_{i-1}$, while $w_{i}=w_{S-\{s_{i}\}}$ conjugates $I_{i-1}$ to $I_{i}$, and so
$I_{i-1}^{u_{i}}=(I_{i-1}^{w^{\prime}_{i-1}})^{w_{i}}=I_{i-1}^{w_{i}}=I_{i}.$
@@ 3.2 Eigenvalues and edge–zeta factors
This fact is closely related to the theory of oppositeness graphs, cf. [undef], and generalizes Proposition 7.0.1 of [undefr]. Both proofs, ours and the one in [undef], are ultimately based on Springer’s theorem on the action of central elements in Iwahori–Hecke algebras.
@@
where $\mathcal{R}$ runs over all conjugacy classes of reflections in the Weyl group $W$, and $\chi_{1}$ denotes the corresponding character of $W$ associated with $\chi$. Then the eigenvalues of $(a_{P_{I}wP_{J}}\,a_{P_{J}wP_{I}})_{r}$ are $Q(w_{I})^{-2}\cdot q^{\,f_{\chi}}$, with multiplicity $d_{\chi}n_{\chi}$, as $\chi$ ranges over all irreducible representations of $G$ with $n_{\chi}\neq 0$.
@@
By Springer’s Theorem 9.2.2 in [undefe], the element $a_{BwB}^{2}$ lies in the center of the Hecke algebra $e_{B}\mathbb{C}Ge_{B}$. Since $e_{P_{I}}\in e_{B}\mathbb{C}Ge_{B}$, we have
@@
Second, we change sides (Lemma 6.1.1 of [undefr]): the eigenvalues of $(a_{BwB}^{2})_{r}$ on $\mathbb{C}Ge_{P_{I}}$ coincide with those of $(a_{BwB}^{2})_{l}$ on $e_{P_{I}}\mathbb{C}G$.
@@
Finally, by Springer’s theorem again, the central element $a_{BwB}^{2}\in e_{B}\mathbb{C}Ge_{B}$ acts on $e_{B}M_{\chi}$, and on $e_{P}M_{\chi}\subseteq e_{B}M_{\chi}$, as scalar multiplication by $q^{f_{\chi}}$. Hence the eigenvalues of $(a_{P_{I}wP_{J}}a_{P_{J}wP_{I}})_{r}$ are
@@ 3.3 Summary: proof and refinement of the main theorem
    $Z_{c}\!\left(X_{2}(\mathcal{B})|_{C},u\right)\;=\;\prod_{\lambda\in\mathop{\rm Spec}\nolimits(T_{C})}\frac{1}{1-\lambda\,u^{\,c}}=\;\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,\left(Q(w_{I})^{-2}\,q^{\,f_{\chi}}\right)^{1/d}\,u^{\,c}\bigr)^{\,m_{C,\chi}(\zeta)d_{\chi}}}.$
We have now proved the Main Theorem 1.0.1, in fact in the following refined form:
@@
    $Z_{c}\!\left(X_{2}(\mathcal{B})|_{C},u\right)=\;\prod_{\chi:\,n_{\chi}\neq 0}\;\prod_{\zeta^{d}=1}\frac{1}{\bigl(1-\zeta\,\left(Q(w_{I})^{-2}\,q^{\,f_{\chi}}\right)^{1/d}\,u^{\,c}\bigr)^{\,m_{C,\chi}(\zeta)d_{\chi}}}.$
@@
$Z\!\left(X_{1}(\mathcal{B}),u\right)=Z_{c}\!\left(X_{2}(\mathcal{B}),u\right)\;=\;\prod_{C}Z_{c}\!\left(X_{2}(\mathcal{B})|_{C},u\right),$
@@
By Tits’ Deformation Principle ([undefr, §8.1], [undefe, §7.4,8.1]), under $q\to 1$ we have $\mathcal{H}\to\mathbb{C}W$, $P_{0}\to W_{0}$, and $D(v_{0},v_{0},c)\to a_{W_{0}uW_{0}}=\tfrac{1}{|W_{0}|}\sum_{x\in W_{0}uW_{0}}x$. The module $e_{P_{I}}M_{\chi}$ deforms to $e_{W_{I}}M_{1,\chi}$, where $M_{1,\chi}$ is the $\mathbb{C}W$-module corresponding to $\chi$. Thus $m_{C,\chi}(\zeta)$ equals the multiplicity of $\zeta$ as an eigenvalue of $a_{W_{0}uW_{0}}$ on $e_{W_{0}}M_{1,\chi}$.